# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ 7 Conclusion
DiDC can handle scenarios where control and treatment groups differ significantly, violating the parallel trends assumption of DiD, or when RDD encounters confounding factors at the threshold. By incorporating more information, DiDC eliminates bias in RDD estimates under specific assumptions about the data-generating processes.
However, it is important to give the necessary importance to the identification assumptions, particularly the time-invariance of confounding effects. We propose a test based on stacked RDDs to assess its validity in practic. Additionally, DiDC requires the treatment effect to be independent of confounding policy, though we introduce a possible relaxation for potential interaction effects.
We find that the DiDC method can eliminate bias completely if the shape of the functions remains stable over time at each side of the threshold. This suggests that it could offer significant advantages over standard RDD estimations, even in settings where no other confounding variables exist at the threshold. We also propose test to compare the derivatives of estimated functions on either side of the threshold, allowing researchers to evaluate the stability of data-generating processes over time.
Monte Carlo simulations demonstrate DiDC’s potential for improvement over RDD, yielding lower bias and better coverage. It can provide more desirable results by incorporating more data, especially when the functional form exhibits minimal temporal variation. Notably, DiDC is the only viable approach when confounding factors render both RDD and DiD unsuitable. The empirical application highlights the importance of the time-invariance assumption.
Future research directions include the development of alternative estimators that are robust to violations of identification assumptions, as well as the exploration of other confidence interval methods specifically tailored for the DiDC design. Overall, the DiDC method offers a valuable addition to the causal inference toolkit. It is applicable in settings where no other methods were previously available and shows potential to reduce bias in estimation in other settings.
@@ Appendix A Setup, assumptions and notation for estimation
We construct the local polynomial estimator following Calonico et al. (2014). For a given $\nu\leq p\in\mathbf{N}$, the general estimand of interest is $\tau_{\nu}=\Delta\mu_{+}-\Delta\mu_{-}$ with $\Delta\mu^{(\nu)}_{+}=\nu!e^{\prime}_{\nu}\delta_{+,p}$, $\Delta\mu^{(\nu)}_{-}=\nu!e^{\prime}_{\nu}\delta_{-,p}$ being the $\nu$th-order derivatives of the $p$th-order local polynomial of the difference. The $p$th-order local polynomial estimators of the $\nu$th-order derivatives $\Delta\mu^{(\nu)}_{+,p}$ and $\Delta\mu^{(\nu)}_{-,p}$ are:
@@
where $e_{\nu}$ is a conformable $(\nu+1)$ unit vector, $K_{h}(u)=K(u/h)/h$, $h_{n}$ is a positive bandwidth sequence, $r_{p}(x)=\begin{bmatrix}1&x&\ldots&x^{p}\end{bmatrix}^{\prime}$, $\Delta Y=\begin{bmatrix}\Delta Y_{1}&\Delta Y_{2}&\ldots&\Delta Y_{n}\end{bmatrix}^{\prime}$. We set $\chi_{n}=\begin{bmatrix}Z_{1}&\ldots&Z_{n}\end{bmatrix}^{\prime}$, $\varepsilon_{\Delta Y}=\begin{bmatrix}\varepsilon_{\Delta Y,1}&\ldots&\varepsilon_{\Delta Y,n}\end{bmatrix}^{\prime}$ with $\varepsilon_{\Delta Y,i}=\Delta Y_{i}-\mu_{\Delta Y}(Z_{i})$, $\mu_{\Delta Y}(Z)=E\left(\Delta Y|Z\right)$ and $\sigma^{2}_{\Delta W,\Delta Y}(Z)=Cov\left(\Delta W,\Delta Y|Z\right)$. We also set
@@
$\Sigma_{\Delta W\Delta Y}$ $=diag\left(\sigma_{\Delta W\Delta Y}^{2}(Z_{i}),\ldots,\sigma_{\Delta W\Delta Y}^{2}(Z_{n})\right)$
@@
$\sigma_{WX+}^{2}=\lim_{z\rightarrow 0^{+}}\sigma_{WX}^{2}(z),\quad\sigma_{WX_{-}}^{2}=\lim_{z\rightarrow 0^{-}}\sigma_{WX}^{2}(z),$
$\sigma_{WX}^{2}(z)=\mathbb{C}[W,X\mid X=z].$
@@
    $\Delta\mu_{-}(z)=E\left[\Delta Y_{i}(0)|Z_{i}=z\right]$ and $\Delta\mu_{+}(z)=E\left[\Delta Y_{i}(1)|Z_{i}=z\right]$ are S times continuously differentiable.
@@ Appendix B Preliminary lemmas and results
Before proceeding, please refer to Appendix A for notation. This appendix restates, with minor adaptations, several lemmas, results and proofs from Calonico et al. (2014) that are necessary for deriving the asymptotic results.
@@
Suppose Assumptions $1-2$ hold, and $nh_{n}\rightarrow\infty$.
@@
For part (a.5), change of variable implies
@@
Proofs for the other items follow similarly to the one above and can be found in Calonico et al. (2014), as they are identical to those provided there. ∎
@@
Suppose Assumptions 1 and 2 hold with $S\geq\ell+2$, and $nh_{n}\rightarrow\infty$.
@@
For part (B), a Taylor series expansion gives
@@
×$\times W_{+}\left(h_{n}\right)Z_{\ell}\left(h_{n}\right)\Gamma_{+,\ell}^{-1}\left(h_{n}\right)H_{\ell}\left(h_{n}\right)e_{s}/n$
@@
×$\times W_{+}\left(h_{n}\right)Z_{\ell}\left(h_{n}\right)\Gamma_{+,\ell}^{-1}\left(h_{n}\right)e_{s}/n$
@@
Suppose Assumptions 1 and 2 hold with $S\geq q+1$, and $n\min\left\{h_{n},b_{n}\right\}\rightarrow\infty$. (B) If $\max\left\{h_{n},b_{n}\right\}\rightarrow 0$, then
@@
×$\times W_{+}\left(b_{n}\right)X_{q}\left(b_{n}\right)\Gamma_{+,q}^{-1}\left(b_{n}\right)e_{p+1}b_{n}^{-p-1}/n^{2}$
@@
Proof for the control group (subindex "-") is analogous. ∎
@@
Suppose Assumptions $1-2$ hold with $S\geq p+2$, and $nh_{n}\rightarrow\infty$. Let $r\in\mathbb{N}$.
@@ Appendix C Proofs of main results
Proof of Claim 1.
Given Assumptions 1 and 2, we can prove that $\tau^{DiDC}=\tau_{1}^{RDD}-\tau_{0}^{RDD}=\Delta Y^{+}-\Delta Y^{-}$.
@@
    Difference of RDDs. From 1 and 2 we can derive two RDDs estimands, one for $t=0$ and one for $t=1$.
    $\tau^{RDD}_{1}$ $=Y_{1}^{+}-Y_{1}^{-}$
    $=E\left[Y_{i,1}\left(1,1\right)-Y_{i,1}\left(0,1\right)-Y_{i,1}\left(1,0\right)+Y_{i,1}\left(0,0\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}D_{1}^{+}-D_{0}^{-}D_{1}^{-}\right)+$
    $\quad\quad E\left[Y_{i,1}\left(0,1\right)-Y_{i,1}\left(0,0\right)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)+$ (C.1)
    $\quad\quad E\left[Y_{i,0}\left(1,0\right)-Y_{i,1}\left(0,0\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}-D_{0}^{-}\right)$
    $\tau^{RDD}_{0}$ $=Y_{0}^{+}-Y0^{-}$
    $=E\left[Y_{i,0}\left(1,0\right)-Y_{i,0}\left(0,0\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}-D_{0}^{-}\right)$ (C.2)
    where $X^{+}=\lim_{\epsilon\rightarrow 0}E(X_{i}|Z_{i}=z_{0}+\epsilon)$ and $X^{-}=\lim_{\epsilon\rightarrow 0}E(X_{i}|Z_{i}=z_{0}-\epsilon)$. By taking the difference between the equations C.1 and C.2 and under assumptions 1-4, we arrive at the differences-in-discontinuity estimand:
@@
    $=E\left[\left(Y_{i,1}(1,1)-Y_{i,1}(0,1)\right)\left(Y_{i,1}(1,0)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}D_{1}^{+}-D_{0}^{-}D_{1}^{-}\right)+$
    $\quad E\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)+$
    $\quad E\left[\left(Y_{i,1}(1,0)-Y_{i,1}(0,0)\right)\left(Y_{i,0}(1,0)-Y_{i,0}(0,0)\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}-D_{0}^{-}\right)$
@@
    RDD of the differences. Using equation3 and taking the difference of the limits above and below the threshold:
@@
    $=E\left[\left(Y_{i,1}(1,1)-Y_{i,1}(0,1)\right)\left(Y_{i,1}(1,0)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}D_{1}^{+}-D_{0}^{-}D_{1}^{-}\right)+$
    $\quad E\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)+$
    $\quad E\left[\left(Y_{i,1}(1,0)-Y_{i,1}(0,0)\right)\left(Y_{i,0}(1,0)-Y_{i,0}(0,0)\right)|Z_{i}=z_{0}\right]\left(D_{0}^{+}-D_{0}^{-}\right)$
@@
Therefore $\tau^{DiDC}=\tau^{RDD}_{1}-\tau^{RDD}_{0}=\Delta Y^{+}-\Delta Y^{-}$. ∎
Proof.
Proof of Lemma 1
Recall the definition $MSE_{\nu,p,s}(h_{n})=E\left[\left(\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}\right)^{2}|\chi_{n}\right]$. We can rewrite as
$MSE_{\nu,p,s}(h_{n})=$ $V\left[\hat{\tau}_{\nu,p,s}(h_{n})|\chi_{n}\right]+\left(E\left[\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}|\chi_{n}\right]\right)^{2}$
@@
$E\left[\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}|\chi_{n}\right]=h_{n}^{1+p-\nu}\frac{\Delta\mu^{(p+1)}_{+}-(-1)^{\nu+p+s}\Delta\mu^{(p+1)}_{-}}{(p+1)!}\nu!e^{\prime}_{\nu}\Gamma_{p}^{-1}\varphi_{p,p+1}\{1+o_{p}(1)\}$ (C.3)
@@
$h_{n}^{MSE}{}_{\nu,p}$ $=arg\min_{h_{n}}MSE_{\nu,p,s}(h_{n})$
$=\left(\frac{\left(1+2\nu\right)\mathbf{V}_{\nu,p}}{n2\left(1+p-\nu\right)\textbf{B}^{2}_{\nu,p,p+1,s}}\right)^{\frac{1}{2p+3}}$
Proof.
Proof of Claim 2
@@
where $\Delta\mu(z)=E\left(\Delta Y_{i}|Z_{i}=z\right)=E\left(Y_{i,1}-Y_{i,0}|Z_{i}=z\right)$.
The notation for the RDD at time $t$ where $\mu_{t}(z)=E\left(Y_{i,t}|Z_{i}=z\right)$ is
@@
It follows that $B\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})\right]=B\left[\hat{\tau}^{RDD}_{1,\nu,p}(h_{n})\right]-B\left[\hat{\tau}^{RDD}_{0,\nu,p}(h_{n})\right]$
Proof.
Proof of Calim 3
$\Delta\mu(z)=$ $E\left[\Delta Y_{i}|Z_{i}=z\right]$
$=$ $E\left[Y_{i,1}-Y_{i,0}|Z_{i}=z\right]$
$=$ $E\left[Y_{i,1}|Z_{i}=z\right]-E\left[Y_{i,0}|Z_{i}=z\right]$

### v2

@@ 8 Conclusion
DiDC can handle scenarios in which the control and treatment groups differ significantly, violating the parallel trends assumption of DiD, or when RDD encounters confounding factors at the threshold. By incorporating more information, DiDC eliminates bias in RDD estimates under specific assumptions about the data-generating processes.
However, it is important to give due attention to the identification assumptions, particularly the time-invariance of confounding effects. We propose a test based on stacked RDDs to assess its validity in practice. Additionally, DiDC requires the treatment effect to be independent of confounding policy, though we introduce a possible relaxation for potential interaction effects.
We find that the DiDC method can eliminate bias completely if the function shapes remain stable on both sides of the threshold over time. This suggests it could offer significant advantages over standard RDD estimators, even in settings where no other confounding variables are present at the threshold. We also propose a test to compare the derivatives of estimated functions on either side of the threshold, allowing researchers to evaluate the stability of data-generating processes over time.
Monte Carlo simulations demonstrate DiDC’s potential to improve upon RDD, yielding lower bias and better coverage. It can provide more desirable results by incorporating more data, especially when the functional form exhibits minimal temporal variation. Notably, DiDC is the only viable approach when confounding factors render both RDD and DiD unsuitable. The empirical application highlights the importance of the time-invariance assumption.
Future research directions include developing robust alternative estimators that are robust to violations of identification assumptions, as well as exploring other confidence interval methods tailored to the DiDC design. Overall, the DiDC method offers a valuable addition to the causal inference toolkit. It is applicable in settings where no other methods were previously available and shows potential to reduce bias in estimation in other settings.
@@ Appendix A Setup, assumptions and notation for estimation
[…]
@@
where $e_{\nu}$ is a conformable $(\nu+1)$ unit vector, $K_{h}(u)=K(u/h)/h$, $h_{n}$ is a positive bandwidth sequence, $r_{p}(x)=\begin{bmatrix}1&x&\ldots&x^{p}\end{bmatrix}^{\prime}$, $\Delta Y=\begin{bmatrix}\Delta Y_{1}&\Delta Y_{2}&\ldots&\Delta Y_{n}\end{bmatrix}^{\prime}$. We define $\chi_{n}=\begin{bmatrix}Z_{1}&\ldots&Z_{n}\end{bmatrix}^{\prime}$, $\varepsilon_{\Delta Y}=\begin{bmatrix}\varepsilon_{\Delta Y,1}&\ldots&\varepsilon_{\Delta Y,n}\end{bmatrix}^{\prime}$ with $\varepsilon_{\Delta Y,i}=\Delta Y_{i}-\mu_{\Delta Y}(Z_{i})$, $\mu_{\Delta Y}(Z)=\mathbb{E}\left(\Delta Y|Z\right)$ and
@@
    $\Delta\mu_{-}(z)=\mathbb{E}\left[\Delta Y_{i}(0)|Z_{i}=z\right]$ and $\Delta\mu_{+}(z)=\mathbb{E}\left[\Delta Y_{i}(1)|Z_{i}=z\right]$ are S times continuously differentiable.
@@ Appendix B Preliminary lemmas and results
Before proceeding, please refer to Appendix A for notation. This appendix restates, with minor adaptations, several lemmas, results and proofs from Calonico et al. (2014b) that are necessary for deriving the asymptotic results.
@@
Suppose Assumptions 121-2 hold, and $nh_{n}\rightarrow\infty$.
@@
For part (a.5), the change of variable implies
@@
Proofs for the other items follow similarly to the one above and can be found in Calonico et al. (2014b), as they are identical to those provided there. ∎
@@
Suppose Assumptions 1 and 3 hold with $S\geq\ell+2$, and $nh_{n}\rightarrow\infty$.
@@
For part (B), a Taylor series expansion yields
@@
$\times W_{+}\left(h_{n}\right)Z_{\ell}\left(h_{n}\right)\Gamma_{+,\ell}^{-1}\left(h_{n}\right)H_{\ell}\left(h_{n}\right)e_{s}/n$
@@
$\times W_{+}\left(h_{n}\right)Z_{\ell}\left(h_{n}\right)\Gamma_{+,\ell}^{-1}\left(h_{n}\right)e_{s}/n$
@@
Suppose Assumptions 1 and 3 hold with $S\geq q+1$, and $n\min\left\{h_{n},b_{n}\right\}\rightarrow\infty$. (B) If $\max\left\{h_{n},b_{n}\right\}\rightarrow 0$, then
@@
$\times W_{+}\left(b_{n}\right)X_{q}\left(b_{n}\right)\Gamma_{+,q}^{-1}\left(b_{n}\right)e_{p+1}b_{n}^{-p-1}/n^{2}$
@@
The proof for the control group is analogous. ∎
@@
Suppose Assumptions 121-2 hold with $S\geq p+2$, and $nh_{n}\rightarrow\infty$. Let $r\in\mathbb{N}$.
@@ Appendix C Proofs of main results
Proof of Lemma 1.
Given Assumptions 1, 2 and 3, we can prove that $\tau^{DiDC}=\tau_{1}^{RD}-\tau_{0}^{RD}=\Delta Y^{+}-\Delta Y^{-}$.
@@
    Difference of RDs. From 2.1 and 2 we can derive two RDs estimands, one for $t=0$ and one for $t=1$.
    $\tau^{RD}_{1}$ $=Y_{1}^{+}-Y_{1}^{-}$
    $=\mathbb{E}\left[Y_{i,1}\left(1,1\right)-Y_{i,1}\left(0,0\right)|Z_{i}=z_{0}\right]$ (C.1)
    $\tau^{RD}_{0}$ $=Y_{0}^{+}-Y0^{-}$
    $=\mathbb{E}\left[Y_{i,0}\left(1,0\right)-Y_{i,0}\left(0,0\right)|Z_{i}=z_{0}\right]$ (C.2)
    where $X^{+}=\lim_{\epsilon\rightarrow 0}E(X_{i}|Z_{i}=z_{0}+\epsilon)$ and $X^{-}=\lim_{\epsilon\rightarrow 0}E(X_{i}|Z_{i}=z_{0}-\epsilon)$. By taking the difference between the equations C.1 and C.2 and adding assumption 4, we arrive at the differences-in-discontinuity estimand:
@@
    $=\mathbb{E}\left[\left(Y_{i,1}(1,1)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]$-
    $\quad\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]$
    $=\mathbb{E}\left[\left(Y_{i,1}(1,1)-Y_{i,1}(1,0)\right)|Z_{i}=z_{0}\right]$
@@
    RD of the differences. Taking the difference of the limits of the first-differenced outcomes above and below the threshold, and under assumptions 1, 2 and 3:
@@
[…]
    $\quad\mathbb{E}\left[\left(Y_{i,1}(0,0)-Y_{i,0}(0,0)\right)|Z_{i}=z_{0}\right]$
    $=\mathbb{E}\left[\left(Y_{i,1}(1,1)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]$-
    $\quad\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]$
@@
Therefore $\tau^{DiDC}=\tau^{RD}_{1}-\tau^{RD}_{0}=\Delta Y^{+}-\Delta Y^{-}$. ∎
Proof of Lemma 3.
Using the derivations in the proof of Lemma 1:
$\tau^{DiDC}$ $=\mathbb{E}\left[\left(Y_{i,1}(1,1)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]$-
$\quad\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]$
Under Assumption 4’,
$\tau^{DiDC}$ $=\mathbb{E}\left[\left(Y_{i,1}(0,1)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]\mathbb{E}\left[\left(Y_{i,1}(1,0)-Y_{i,1}(0,0)\right)|Z_{i}=z_{0}\right]$-
$\quad\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]$
By Assumption 4, the second factor is time-invariant and can be replaced with its period-0 analogue, which is exactly $\tau^{RD}_{0}$. Substituting this expression gives
$\tau^{DiDC}=(\mathbb{E}\!\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)\mid Z_{i}=z_{0}\right]+1)\tau^{RD}_{0}$
Rearranging yields the desired result,
$\frac{\tau^{DiDC}}{\tau^{RD}_{0}}+1=\mathbb{E}\!\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)\mid Z_{i}=z_{0}\right].$
Proof of Lemma 4.
Recall the definition $MSE_{\nu,p,s}(h_{n})=\mathbb{E}\left[\left(\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}\right)^{2}|\chi_{n}\right]$. We can rewrite as
$MSE_{\nu,p,s}(h_{n})=$ $V\left[\hat{\tau}_{\nu,p,s}(h_{n})|\chi_{n}\right]+\left(\mathbb{E}\left[\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}|\chi_{n}\right]\right)^{2}$
@@
$\mathbb{E}\left[\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}|\chi_{n}\right]=h_{n}^{1+p-\nu}\frac{\Delta\mu^{(p+1)}_{+}-(-1)^{\nu+p+s}\Delta\mu^{(p+1)}_{-}}{(p+1)!}\nu!e^{\prime}_{\nu}\Gamma_{p}^{-1}\varphi_{p,p+1}\{1+o_{p}(1)\}$ (C.3)
@@
$h_{n,\nu,p}^{MSE}$ $=arg\min_{h_{n}}MSE_{\nu,p,s}(h_{n})$
$=\left(\frac{\left(1+2\nu\right)\mathbf{V}_{\nu,p}}{2n\left(1+p-\nu\right)\textbf{B}^{2}_{\nu,p,p+1,s}}\right)^{\frac{1}{2p+3}}$
Proof of Lemma 4.
@@
where $\Delta\mu(z)=\mathbb{E}\left(\Delta Y_{i}|Z_{i}=z\right)=\mathbb{E}\left(Y_{i,1}-Y_{i,0}|Z_{i}=z\right)$.
The notation for the RD at time $t$ where $\mu_{t}(z)=\mathbb{E}\left(Y_{i,t}|Z_{i}=z\right)$ is
@@
It follows that $B\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})\right]=B\left[\hat{\tau}^{RD}_{1,\nu,p}(h_{n})\right]-B\left[\hat{\tau}^{RD}_{0,\nu,p}(h_{n})\right]$
Note that
$\Delta\mu(z)=$ $\mathbb{E}\left[\Delta Y_{i}|Z_{i}=z\right]$
$=$ $\mathbb{E}\left[Y_{i,1}-Y_{i,0}|Z_{i}=z\right]$
$=$ $\mathbb{E}\left[Y_{i,1}|Z_{i}=z\right]-\mathbb{E}\left[Y_{i,0}|Z_{i}=z\right]$